Consider the following premerger information about a bidding firm (Firm A) and a target firm (Firm B). Assume that both firms have no debt outstanding.

Firm A

Firm B

Share price

50

20

Number of shares

10,000

3,000

Firm A has estimated that the value of the synergistic benefits from acquiring Firm B is $50,000.

If Firm B is acquired for $30 per share in cash, what is the merger premium in this merger?

If Firm B is acquired for $30 per share in cash, what is the NPV of the merger for Firm A?

What will the price per share of the merged firm be assuming the conditions in the previous question?
Answer #1

Ans 1: Merger premium=Value of Firm B+Synergetic benefits from acquiring- Cash paid for merger

=60000+50000-90000

   =20000

Ans 2:NPV of merger=Value of Firm A+ Value of Firm B+ Synergetic Benefits- Cash paid for merger

=500000+60000+50000-90000

   =520000

Ans 3:Price per share of merged firm= (Market value of Firm A+ Market value of Firm B+ Synergetic benefits)\divTotal no. of shares in the merged firm

For this we need to find the no. of shares of the merged firm

No. of shares of Firm A=10000

No. of shares of Firm A given to Firm B=Purchase Price\divMarket price per share of Firm B

   =90000\div50

   =1800

So total no.of shares is 11800

Price per share of merged firm=(500000+60000+50000)\div11800

   =51.69